solve each equation by factoring..
1. 4x2 – 17x – 15 = 0
2. 10x2 + 19x + 6 = 0
3. Solve 3x2 = 4,800 by using square roots.

Answer by scott8148(6628)

1. factors of 4 and -15 that "combine" to give -17 __ (4x+3)(x-5)=0

2. factors of 10 and 6 that "combine" to give 19 __ (5x+2)(2x+3)=0



3. dividing by 3 __ x^2=1600

taking square root __ x=±40